OpenStudy (anonymous):

Keith is an independent filmmaker. He's planning to serve pizza and sub sandwiches at his first film debut. The number of sandwiches is twice the value of four less than the number of pizzas. If he ordered 62 sandwiches, how many pizzas did he order? 48 35 33 12

OpenStudy (anonymous):

Pizza = x Sandwiches = 2(x-4) So if sandwiches = 62 62= 2(x-4), multiply 2 into the parathesis 62 = 2x - 8 (add 8 to both sides) 70 = 2x (divide both sides by 2) x = 35